OpenStudy (anonymous):

@Godlovesme Which shows three ordered pairs that are all solutions of the equation y = x + 4? A. (1, 4), (2, 6), and (3, 7) B. (1, 4), (2, 8), and (3, 12) C. (3, 7), (5, 9), and (7, 10) D. (3, 7), (5, 9), and (7, 11)

Mehek (mehek14):

plug in the ordered pairs

OpenStudy (anonymous):

?

Mehek (mehek14):

lets try the first one in A 4 = 1 + 4 is that correct?

Mehek (mehek14):

does 4 = to 4 + 1?

OpenStudy (anonymous):

how did you already get 2 medals

Mehek (mehek14):

4+1 = 5

OpenStudy (anonymous):

No sorry

Mehek (mehek14):

ok so since one of he answers in option A is wrong, then A is not the answer

OpenStudy (anonymous):

Ok

Mehek (mehek14):

lets move to B (1,4) 4 = 1 + 4? This is the same like the first one' is it correct?

OpenStudy (anonymous):

Nope

Mehek (mehek14):

ok so A and B can't be the answer

Mehek (mehek14):

Now C (3,7) 7 = 3 + 4?

OpenStudy (anonymous):

Yes

Mehek (mehek14):

ok lets try the second one in C (5,9) 9 = 5 + 4?

OpenStudy (anonymous):

Yes

Mehek (mehek14):

ok now the third one in C (7,10) 10 = 7 + 4?

OpenStudy (anonymous):

No

Mehek (mehek14):

so it can't be C

Mehek (mehek14):

That leaves you with D

OpenStudy (anonymous):

Thank you =D can you help me with more?

Mehek (mehek14):

Sure but can you open another post since I'm going to have to scroll up and down a lot?

OpenStudy (anonymous):

Sure